Can you help me with the below problem? Trying to get the steps down before I move on to more difficult problems.
Find the equation of the line, in slope intercept form, which passes through the points (-1,-2) and (4,8).
Thank you.

Answer by stanbon(48558)

Find the equation of the line, in slope intercept form, which passes through the points (-1,-2) and (4,8).
------------------
You need the slope and the y-intercept.
slope=[8--2]/[4--1]= 10/5=2
y=mx+b
8=2(4)+b
b=0
EQUATION:
y=2x
Cheers,
Stan H.